OpenStudy (anonymous):

Can someone help me with 19= h/3 -8 I don't understand

OpenStudy (sloppycanada):

I believe you're trying to solve for "h" right?

OpenStudy (sloppycanada):



OpenStudy (anonymous):

yes

OpenStudy (sloppycanada):

Ah okay, the first step you want to do is multiply both sides by (3-8). This is because we want "h" by himself.

OpenStudy (anonymous):

\[19=\frac{ h }{ 3-8 }\]\[19=\frac{ h }{ -5 }\]Multiply -5 on both sides

OpenStudy (sloppycanada):

I suppose that works too.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

thanks

OpenStudy (sloppycanada):

19 x -5?